70 likes | 142 Views
We identified a noise issue in the -6V of the pre-amp but its source is unknown. Using a different low voltage setting, we measured pkpk after the pre-amp with 10,000 events, obtaining 1.279 ± 0.030V through Gaussian fitting. Next, we examined the Ref. PMT from the USTC module, recording values of 690.7 ± 13.3 (1.9%) and 718.9 ± 17.3 (2.4%). Despite the IHEP DAQ program not completing, we utilized the USTC program to acquire data, though limited to reading 88 TOF data due to bugs. Analyzing T(sub) = T(tof) - T(ref) with Gaussian fitting revealed inconsistencies in laser signal distribution. Investigation continues as certain east TOF sections show signal while west TOF does not. Possible reasons include insufficient HV. In conclusion, system functionality is confirmed. Further tests are planned to verify fiber integrity. Additional steps involve the installation of west 17 bundles, with 7 remaining. - LIU Qian
E N D
12.20 report LIU Qian
We found a noise on the -6V of pre-amp. We don’t know where it comes from. 50 ms/div DC1M 2.00 V/div
So we used another low voltage, and measured pkpk after pre-amp. 10,000 events. Fit with the histogram with Gaussian, we got: 1.279 ± 0.030 V. reflection 20 ns/div
Then read Ref PMT from USTC module 0.03 pC/channel 690.7 ± 13.3 (1.9%) 718.9 ± 17.3 (2.4%)
After that, we used the USTC module to measure the laser signal. The IHEP DAQ program still doesn’t finish. So we have to use the USTC program to acquire the data. (but USTC program has some bugs, and can only read 88 TOF data). We set all HV as 1050V, closed the west shuttle. Then we analysis the T(sub) = T(tof) – T(ref), and fit with Gaussian. Technically, only the west TOF end can have the signal because we only pulsed the east side laser, the east TOF end can only have some reflection laser. But we found that some east TOF does have signal, and west TOF doesn’t have signal. We’re still looking for the reason. Maybe the HV is not enough.
The mean and sigma of west end TOF T(sub) We’re looking for the reason.
Conclusion • At least, everything is working now. • I do really want to check all fibers are OK first, but I think the DAQ won’t be ready. Maybe I should check them one by one by the oscilloscope. • We installed west 17 bundles to the distribution boxes today. There are only 7 bundles left now. LIU Qian